# Generative Flows on Synthetic Pathway for Drug Design

### v1

@@ B.1 Forward policy
For the model architecture, we used a graph transformer (Yun et al., 2022) as the backbone $f_{\theta}$ following Bengio et al. (2021) and a multi-layer perceptron (MLP) for action embedding $g_{\theta}$. The graph embedding dimension is $d_{1}$, and the building block embedding dimension is $d_{2}$. The molecular graph for a state is $s$, and the GFlowNet condition vector is $c$ which includes a reward exponent and multi-objective optimization weights (Jain et al., 2023). $a\|b$ means the concatenation of two feature vectors $a$ and $b$.
@@
B.4 Non-hierarchical Markov decision process
Table 6: Comparison with SynFlowNet and RGFN. RGFN investigates scaling up to 64,000 building blocks, but their experimental validation and proof-of-principle implementations use only 350.
@@ B.4 Non-hierarchical Markov decision process
While RxnFlow shares similar rules with SynFlowNet (Cretu et al., 2024) and RGFN (Koziarski et al., 2024), there are several major differences as described in Table 6. In particular, we formulate the non-hierarchical MDP instead of the hierarchical MDP for bi-molecular reactions.
@@ C.3 Experimental details
Pocket-specific optimization.
For pocket-specific optimization with GPU-accelerated UniDock, we normalize the Vina and SA scores for multi-objective optimization:
@@
Pocket-conditional generation in a zero-shot manner.
We used the modified version of the TacoGFN’s reward function and training set. Since we don’t need to optimize SA (Ertl & Schuffenhauer, 2009), we excluded the SA term from the reward functions:
@@
$\text{Reward}(x)=\frac{r_{\text{affinity}}(x)\times r_{\text{QED}}(x)}{\sqrt[3]{\text{HeavyAtomCounts}(x)}}$ (31)
@@
Introducing further objectives without retraining.
For the restricted block library (TPSA$<$30), we set the action space subsampling ratio as 10% for both AddFirstReactant and ReactBi, and we set that as 1% for an entire library.
Scaling action space without retraining.
For action space subsampling, we set the subsampling ratio as 2% for both AddFirstReactant and ReactBi for the “seen” and “unseen” libraries. For the “all” library (“seen” + “unseen”), we set the subsampling ratio as 1%.
Ablation study.
We set different subsampling ratios according to the building block library size. For 100-sized, 1k-sized, and 10k-sized libraries, we do not perform the action space subsampling. We set an action space subsampling ratio of 10% for a 100k-sized one and 1% for a 1M-sized one.
@@ C.4 Softwares
Molecular docking software.
For a fair comparison with the baseline model, we used UniDock (Yu et al., 2023) for target-specific generation and QuickVina 2.1 (Alhossary et al., 2015) for SBDD. The initial ligand conformer is generated with srETKDG3 (Wang et al., 2020) in RDKit (Landrum et al., 2013). For QuickVina, we converted the molecule format to pdbqt with OpenBabel (O’Boyle et al., 2011) and AutoDock Tools (Huey et al., 2012). To set up an exhaustive search, we set the search mode to balance for UniDock and the exhaustiveness to 8 for QuickVina. We kept the seed fixed at 1 throughout the ETKDG and the whole docking process.
Docking proxy.
We used the QuickVina proxy proposed by Shen et al. (2023) which is implemented in PharmacoNet (Seo & Kim, 2023). We used a proxy model trained on the CrossDocked2020 training set rather than the model trained on the ZINCDock15M training set.
Synthetic accessibility estimation.
To evaluate the synthetic accessibility of molecules, we used the retrosynthesis planning tool AiZynthFinder (Genheden et al., 2020). AiZynthFinder uses MCTS to find synthesis paths and estimate the number of steps, search time, success rate, and synthetically accessible score as metrics to indicate synthesis complexity or synthesizability.
@@ C.5 Baselines
SynNet.
For SynNet (Gao et al., 2022b), we perform multi-objective optimization with the following reward function:
@@
RGFN.
Since the code of RGFN is not released, we reimplement the RGFN.
BBAR.
Since BBAR (Seo et al., 2023) allows multi-conditional generation, we directly used QED and docking scores without any processing. We split 64,000 ZINC20 molecules according to the reported splitting of BBAR: 90% for the training set, 8% for the validation set, and 2% for the test set (in our case, the number of sampling molecules). We performed UniDock for training and validation set to prepare the label of the molecules. Since BBAR requires the desired property value, we used the average docking score of the top 100 diverse modes from our model.
Pocket2Mol, TargetDiff, DecompDiff, TacoGFN.
We followed the reported generative setting to generate 100 molecules for each CrossDocked test pocket. We set the center of the pockets with the reference ligands in the CrossDocked2020 database. We reuse reported runtime in Shen et al. (2023), which is measured on NVIDIA A100 for Pocket2Mol, TargetDiff, and DecompDiff, and NVIDIA RTX3090 for TacoGFN.
@@ D.1 Additional results for Pocket-specific generation task
Percentage of Synthesizable Molecules (%, $\uparrow$)
@@
D.2 Additional results for scaling action space without retraining
@@
D.3 Theoretical analysis
@@
As shown in Figures 10(a) and 10(b), we compare a baseline GFlowNet trained without subsampling (“base”) to models using various subsampling ratios and Monte Carlo (MC) sampling. The differences in $\log Z_{\theta}$ are relatively small ($<$0.005) across all settings, and increasing MC samples for state flow estimation $F_{\theta}$ further reduced the bias. In Figures 10(c) and 10(d), we also evaluate the bias in the trajectory balance loss (^$\hat{\mathcal{L}}_{\text{TB}}$) and its gradient norm ($\|\nabla_{\theta}\hat{\mathcal{L}}_{\text{TB}}\|$) during training, finding negligible differences compared to the true values. These results indicate that our importance sampling reweighting approach effectively mitigates bias from action space subsampling, enabling efficient and accurate policy estimation.

### v2

@@ 5 Conclusion
Acknowledgement
This work was supported by Basic Science Research Programs through the National Research Foundation of Korea (NRF), grant-funded by the Ministry of Science and ICT (NRF-2023R1A2C2004376, NRF-2022M3J6A1063021, RS-2023-00257479). In addition, this research was supported by a grant of Korean ARPA-H Project through the Korea Health Industry Development Institute (KHIDI), funded by the Ministry of Health & Welfare, Republic of Korea (RS-2024-00512498).
@@ B.1 Forward policy
We use the model architecture inspired from Cretu et al. (2024); Koziarski et al. (2024). We used a graph transformer (Yun et al., 2022) as the backbone $f_{\theta}$ following Bengio et al. (2021) and a multi-layer perceptron (MLP) for action embedding $g_{\theta}$. The graph embedding dimension is $d_{1}$, and the building block embedding dimension is $d_{2}$. The molecular graph for a state is $s$, and the GFlowNet condition vector is $c$ which includes a reward exponent and multi-objective optimization weights (Jain et al., 2023). $a\|b$ means the concatenation of two feature vectors $a$ and $b$.
@@
3D interaction modeling.
Instead of using a 2D molecular graph of the ligand, the model can utilize a 3D binding complex graph to represent the state as the input of the graph transformer backbone $f_{\theta}$. In this graph, each ligand atom connects to its $K_{1}$ nearest protein atoms as well as to its neighboring ligand atoms, while each protein atom connects to its $K_{2}$ nearest protein atoms. The edges encode spatial relationship information between the connected nodes.
Furthermore, we lightweight the MDP formulation to implement computationally intensive 3D interaction modeling with modified layers $\text{MLP}^{\ast}_{\texttt{AddFirstReactant}}:\mathbb{R}^{d_{1}}\rightarrow\mathbb{R}^{d_{2}}$ and $\text{MLP}^{\ast}_{\texttt{ReactBi}}:\mathbb{R}^{d_{1}+|\mathcal{R}_{2}|}\rightarrow\mathbb{R}^{d_{2}}$ as follows:
$F_{\theta}(s_{0},b,c)=\text{MLP}^{\ast}_{\texttt{AddFirstReactant}}(f_{\theta}(s_{0},c))\odot g_{\theta}(b),$ (29)
$F_{\theta}(s,(r_{2},b),c)=\text{MLP}^{\ast}_{\texttt{ReactBi}}(f_{\theta}(s,c)\|\delta(r_{2}))\odot g_{\theta}(b),$ (30)
@@
B.4 GFlowNet Training.
The training algorithm with the trajectory balance objective is described at Algorithm 1. The notations are in Sec. 3.3.
Algorithm 1 Training GFlowNets with action space subsampling
1:input Entire action space $\mathcal{A}$, Maximum trajectory length $N$
2:repeat
6:until model converges
B.5 Comparison between three GFlowNet methods
[…]
@@ B.5 Comparison between three GFlowNet methods
[…]
While RxnFlow shares similar rules with SynFlowNet (Cretu et al., 2024) and RGFN (Koziarski et al., 2024), there are several major differences as described in Table 6. Since there are two versions of SynFlowNet, the version presented at the ICLR 2024 GEM workshop and the version updated on October 16, 2024, we refer to the updated version as SynFlowNet(v2024.10). For the benchmark studies, we use the workshop version of SynFlowNet(v2024.5).
Backward Trajectory. In contrast to the simple fragment-based or atom-based GFlowNet, some backward transitions do not reach the initial state $s_{0}$. While RGFN and SynFlowNet do not consider the invalid transitions, SynFlowNet(upd) and RxnFlow address this issue. To prevent the invalid backward transition, SynFlowNet(upd) introduces additional maximum likelihood and REINFORCE (Williams, 1992) training objectives for parameterized backward policy to prefer the backward transitions which can reach the initial state. In contrast, we implement explicit retrosynthetic analysis within maximum reaction steps for each state to collect valid backward trajectories.
Flow Network Framework. In contrast to SynFlowNet which uses discrete action space, RGFN, SynFlowNet(upd), and RxnFlow include the action embedding (Dulac-Arnold et al., 2015) for chemical building blocks. While every building block must be sampled and trained at least once to prevent unsafe actions in a discrete action space, action embedding expresses actions based on structural information–domain knowledge. In particular, we formulate the non-hierarchical MDP instead of the hierarchical MDP for bi-molecular reactions. It is adaptable to modified building block libraries as described in Sec. B.6 but increases the computational cost and memory requirement dramatically. Thanks to action space subsampling, RxnFlow can use non-hierarchical MDP without a computational burden problem.
Sampling algorithm. For each forward transition step, RGFN and SynFlowNet (both versions) model the flow for the entire action space to estimate the forward transition probability. Therefore, the computational cost and memory requirement is proportional to the number of using building blocks $\mathcal{B}$, i.e., $\mathcal{O}(|\mathcal{B}|)$. Instead of modeling all actions, we propose an action space subsampling technique, which is similar to the negative sampling technique (Mikolov et al., 2013) in natural language processing, to estimate the forward transition probability from the subset of action space. It makes a cost-variance trade-off to decrease the computational cost complexity to $\mathcal{O}(r|\mathcal{B}|)$, where $r$ is a subsampling ratio. As a result, RxnFlow can handle millions of building blocks on training and inference with the computational cost of handling 10,000 building blocks.
B.6 Non-hierarchical Markov decision process
@@ C.3 Experimental details
Pocket-specific optimization. For pocket-specific optimization with GPU-accelerated UniDock, we normalize the Vina and SA scores for multi-objective optimization:
@@
Pocket-conditional generation in a zero-shot manner. We used the modified version of the TacoGFN’s reward function and training set. Since we don’t need to optimize SA (Ertl & Schuffenhauer, 2009), we excluded the SA term from the reward functions:
@@
$r_{\text{SA}}(x)=\begin{cases}\widehat{\text{SA}}(x)/0.8&\text{if}~{}\widehat{\text{SA}}(x)\leq 0.8\\ 1&\text{otherwise}\\ \end{cases}$
[…]
@@
Introducing further objectives without retraining. For the restricted block library (TPSA$<$30), we set the action space subsampling ratio as 10% for both AddFirstReactant and ReactBi, and we set that as 1% for an entire library.
Scaling action space without retraining. For action space subsampling, we set the subsampling ratio as 2% for both AddFirstReactant and ReactBi for the “seen” and “unseen” libraries. For the “all” library (“seen” + “unseen”), we set the subsampling ratio as 1%.
Ablation study. We set different subsampling ratios according to the building block library size. For 100-sized, 1k-sized, and 10k-sized libraries, we do not perform the action space subsampling. We set an action space subsampling ratio of 10% for a 100k-sized one and 1% for a 1M-sized one.
Further improvement with 3D interaction modeling. To better analyze the effects of interaction modeling, we limit the maximum reaction step to 1. Since the initial state is an empty graph, we can analyze the impact of interaction modeling on a single decision process for selecting a reaction to perform. To obtain the 3D binding conformation, we used GPU-accelerated UniDock (Yu et al., 2023) under the fast search mode. We used the same neural network structure for both 2D-based generation and 3D-based generation. For the reward setting, we used the Vina docking score as a reward function and filtered out molecules according to the Lipinski Rule.
@@ C.4 Softwares
Molecular docking software. For a fair comparison with the baseline model, we used UniDock (Yu et al., 2023) for target-specific generation and QuickVina 2.1 (Alhossary et al., 2015) for SBDD. The initial ligand conformer is generated with srETKDG3 (Wang et al., 2020) in RDKit (Landrum et al., 2013). For QuickVina, we converted the molecule format to pdbqt with OpenBabel (O’Boyle et al., 2011) and AutoDock Tools (Huey et al., 2012). To set up an exhaustive search, we set the search mode to balance for UniDock and the exhaustiveness to 8 for QuickVina. We kept the seed fixed at 1 throughout the ETKDG and the whole docking process.
Docking proxy. We used the QuickVina proxy proposed by Shen et al. (2023) which is implemented in PharmacoNet (Seo & Kim, 2024). We used a proxy model trained on the CrossDocked2020 training set rather than the model trained on the ZINCDock15M training set.
Synthetic accessibility estimation. To evaluate the synthetic accessibility of molecules, we used the retrosynthesis planning tool AiZynthFinder (Genheden et al., 2020). AiZynthFinder uses MCTS to find synthesis paths and estimate the number of steps, search time, success rate, and synthetically accessible score as metrics to indicate synthesis complexity or synthesizability.
@@ C.5 Baselines
SynNet. For SynNet (Gao et al., 2022b), we perform multi-objective optimization with the following reward function:
@@
RGFN. Since the code of RGFN is not released, we reimplement the RGFN.
BBAR. Since BBAR (Seo et al., 2023) allows multi-conditional generation, we directly used QED and docking scores without any processing. We split 64,000 ZINC20 molecules according to the reported splitting of BBAR: 90% for the training set, 8% for the validation set, and 2% for the test set (in our case, the number of sampling molecules). We performed UniDock for training and validation set to prepare the label for the molecules. Since BBAR requires the desired property value, we used the average docking score of the top 100 diverse modes from our model.
Pocket2Mol, TargetDiff, DecompDiff, TacoGFN. We followed the reported generative setting to generate 100 molecules for each CrossDocked test pocket. We set the center of the pockets with the reference ligands in the CrossDocked2020 database. We reuse reported runtime in Shen et al. (2023), which is measured on NVIDIA A100 for Pocket2Mol, TargetDiff, and DecompDiff, and NVIDIA RTX3090 for TacoGFN.
DiffSBDD, MolCRAFT. We used the generated samples from their official GitHub repository.
@@ D.1 Additional results for Pocket-specific generation task
AiZynthFinder Success Rate (%, $\uparrow$)
@@
D.2 Property distribution for pocket-specific generation task
We reported the property distribution of the generated molecules for FragGFN, SynFlowNet, RGFN, and RxnFlow for each of 15 LIT-PCBA targets.
Figure 9: The property distribution of the generated samples for the first 10 LIT-PCBA targets.
Figure 10: The property distribution of the generated samples for the last 5 LIT-PCBA targets.
D.3 Scaling laws with baseline GFlowNets
In this section, we investigate the scaling laws of our model and the baseline GFlowNets, SynFlowNet and RGFN, focusing on performance (Figure 11) and computational cost (Figure 12). Our action space subsampling method reduces the computational cost and memory consumption via cost-variance trade-off and memory-variance trade-off.
Figure 11: Optimization power and diversity. Average of standard deviation over the 4 runs. (a) Average docking score. (b) The uniqueness of Bemis-Murcko scaffolds. (c) Average Tanimoto distance.
Figure 12: Runtime according to the building block library size. Average of standard deviation over the 100 batches. (a) The training runtime. (b) The sampling runtime without model training.
Performance.
We conducted an optimization for the kappa-opioid receptor using RxnFlow and baseline GFlowNets. To prevent the hacking of docking by increasing the molecular size, we performed Vina-QED multi-objective using the multi-objective GFlowNet framework (Jain et al., 2023). As shown in Figure 11, all three models exhibit similar trends in docking score optimization. However, conventional GFlowNets face memory resource constraints that scale with the action space size. As a result, SynFlowNet and RGFN were restricted to library sizes up to 100,000 and 500,000 sizes, respectively, due to memory limitations. In contrast, RxnFlow can accommodate larger action spaces leveraging the memory-variance trade-off.
Speed.
While SynFlowNet and RGFN consider all actions in the massive action space to estimate the forward transition probability, our architecture estimate Our architecture estimates forward transition probabilities over a subset of the action space, unlike SynFlowNet and RGFN, which compute all flows for the entire action space. This design allows RxnFlow to handle larger action spaces without encountering computational bottlenecks. To evaluate the cost-efficiency of our technique, we measured runtime during both training and generation. To unify the generation environments across different methods, we used a constant reward function $R(x)=1$ and a maximum reaction step of 1. Training times were measured on random trajectories, and sampling times were measured with initialized models, according to the building block library size. Due to memory limitations, SynFlowNet was restricted to library sizes up to 400,000.
For RxnFlow, we tested two configurations: RxnFlow (1k) and RxnFlow (10k), which sample up to 1,000 and 10,000 building blocks, respectively. As demonstrated in Figure 12, RxnFlow achieves significantly better cost efficiency in both training and generation compared to the baseline models. Additionally, computational costs can be easily reduced by adjusting the action-space subsampling ratio.
D.4 Statistical information for pocket-conditional generation task
Table 13: Statistical Information for pocket-conditional generative models. Mean and standard deviation for 5 sample sets.
Vina ($\downarrow$) QED ($\uparrow$)
Category Model Avg. Std. Avg. Std.
Atom Pocket2Mol -7.603 0.087 0.567 0.007
TargetDiff -7.367 0.028 0.487 0.006
DiffSBDD -6.949 0.079 0.467 0.002
DecompDiff -8.350 0.033 0.368 0.004
MolCRAFT -8.053 0.033 0.500 0.004
MolCRAFT-large -9.302 0.033 0.448 0.002
Fragment TacoGFN -8.237 0.268 0.671 0.002
Reaction RxnFlow -8.851 0.031 0.666 0.001
D.5 Target specificity of generated samples
To investigate the target specificity of pocket-conditional generation, we measured delta score (Gao et al., 2024a) for the top-10 molecules for each pocket. The delta score evaluates the pocket specificity of a proposed molecule by comparing the docking scores difference in how well each molecule binds to other proteins compared to the target protein.
Table 14: Delta Score for each methods
Category Model Delta Score
Atom Pocket2Mol -1.74
Atom DecompDiff -1.29
Atom DiffSBDD -1.02
Atom MolCRAFT -2.08
Fragment TacoGFN -1.13
Reaction RxnFlow -1.13
D.6 Ablation study for non-hierarchical Markov decision process structure
Figure 13: Reward distribution of generated molecules under different building block libraries: a randomly sampled building block (General) and a nonaromatic building block set (Nonaromatic). (a) The non-hierarchical MDP. (b) The hierarchical MDP.
To investigate the effectiveness of the non-hierarchical MDP structure, we performed an ablation study. We randomly selected 100,000 general building blocks from the Enamine building block library for model training (general). To simulate a scenario where specific functional groups are unallowed, we filtered out all aromatic building blocks to create a nonaromatic building block set. We trained the GFlowNets under Vina-QED multi-objective settings (Jain et al., 2023) against the beta-2 adrenergic receptor for 1,000 training oracles. After training, we generated 100 molecules using both the general block set and the nonaromatic block set without additional training.
As shown in Figure 13, the proposed non-hierarchical MDPs closely align with the identified reward distributions for both objectives, Vina and QED, on the general and nonaromatic building block sets. In contrast, hierarchical MDPs, as utilized in existing methodologies, demonstrate a shift in the reward distribution when the building block set is restricted. This indicates that non-hierarchical MDPs are more robust in changes in the building block set compared to hierarchical MDPs.
D.7 Additional results for scaling action space without retraining
@@
Moreover, we investigated the generalization ability of our method to structurally different building blocks under the Vina-QED objectives against the beta-2 adrenergic receptor. For model training, we randomly selected 100,000 building blocks from the entire library (seen). Additionally, we selected 100,000 building blocks with a Tanimoto similarity of less than 0.5 to all training building blocks (unseen). The model was trained using the seen blocks first, and then the trained model subsequently generate molecules with seen blocks, unseen blocks, and a combination of both sets (all), respectively. As illustrated in Figure 15, the model can also generate samples with similar reward distributions from building blocks with different distributions than the ones used for training.
Figure 15: Reward distribution of generated molecules.
D.8 Theoretical analysis
@@
As shown in Figures 17(a) and 17(b), we compare a baseline GFlowNet trained without subsampling (“base”) to models using various subsampling ratios and Monte Carlo (MC) sampling. The differences in $\log Z_{\theta}$ are relatively small ($<$0.005) across all settings, and increasing MC samples for state flow estimation $F_{\theta}$ further reduced the bias. In Figures 17(c) and 17(d), we also evaluate the bias in the trajectory balance loss (^$\hat{\mathcal{L}}_{\text{TB}}$) and its gradient norm ($\|\nabla_{\theta}\hat{\mathcal{L}}_{\text{TB}}\|$) during training, finding negligible differences compared to the true values. These results indicate that our importance sampling reweighting approach effectively mitigates bias from action space subsampling, enabling efficient and accurate policy estimation.